OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = 4(4n+1)(8n+7)/6 prove that n=1

OpenStudy (anonymous):

what are the dots ?

OpenStudy (anonymous):

all you really have to do is plug 1 into all of the n's .

OpenStudy (anonymous):

4 time 6 , mulitply

OpenStudy (campbell_st):

well start by testing for n = 1 and show the 1st term is equal to the sum of 1 terms.... i.e. the left hand side equals the right hand side

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

4(4(1)+1)(8(1)+7)/6

OpenStudy (anonymous):

@campbell_st can walk you through this i am sure i just want to ask if you understand how to do a proof by induction, i.e. what the steps are after understanding what you need to do, there is only a bit of algebra to do it

OpenStudy (campbell_st):

well I think the problem fails for n = 1 so it can't be proven

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

how would the problem be false

OpenStudy (campbell_st):

did you get the sum of 1 term, n = 1, as 50..?

OpenStudy (anonymous):

i have 4x1(4x1+2)=4x(4+2)=4x6

OpenStudy (anonymous):

=24

OpenStudy (campbell_st):

and the 1st term is 4.6 which is 24.. so the sum of 1 term doesn't equal the 1st term...

OpenStudy (campbell_st):

so all you can do is show that the sum of the terms is false as it fails at n = 1

OpenStudy (campbell_st):

@melchar do you understand the how to do induction...?

OpenStudy (anonymous):

not really

OpenStudy (campbell_st):

ok... its a 3 step process step 1: prove the formula is true for the smallest value, which is usually n = 1 so is the 1st term equal to the sum of 1 terms which means you substitute n = 1 into the right hand side of the equation in you question you found the sum of 1 term was 50 while the 1st term was 24. So it couldn't be proven Step 2: Assume that the formula is true for sum integer, n = k so in your question you would have said the sum of k terms is \[s_{k} =\frac {4(4k +1)(8k + 7)}{6}\] Step 3: Prove that the formula is true for the term n = k+1 Which means you need to show the sum to k plus the (k +1) term is equal to the sum of k + 1 terms. \[S_{k +1} = S_{k} + a_{k + 1}\] so you need to use the assumed sum, from step 2 and add the next term This is the algebra part \[\frac{4(4(k + 1) + 1)(8(k + 1)+7)}{6} = \frac{4(4k + 1)(8k + 7)}{6} + 4(k + 1)(4(k + 1 + 2)\] hope that makes some sense

OpenStudy (campbell_st):

here are some notes from Sydney University on doing mathematical induction http://sydney.edu.au/stuserv/documents/maths_learning_centre/induction.pdf

OpenStudy (anonymous):

okay thank you very much !!! really helped